Mutt Inc has 5.5% coupon bonds that have 3.5 years to maturity and a price of $978.56. What is the yield to maturity?

Mutt has another bond with a 4.75% coupon with 6 years to maturity and a yield to maturity of 3.94%. What is the price of the bond?

GW Corp issued a 10-year bond 2 years ago with a coupon rate of 5.8%. If the yield to maturity on the bond is 6.6%, what is the current price?

GW Corp also has just issued an 8-year discount bond for an issue price of $661.56. What is the yield to maturity?
Answer #1

Calculation for each is shown below

в с D E 1 Using the rate function in excel we can calculate YTM for Mutt Inc bond 2 No of payments (NPER) 3.50 3 Price (PV) $

Following shows the working

3.5 978.56 =1000*5.5% 1000 =RATE(B2,-B4,B3,-1000) 1 Using the rate function in excel we can calculate YTM for Mutt Inc bond 2